# Minimal Model for Reservoir Computing

### v2

@@ Abstract
A minimal model for reservoir computing is studied. We demonstrate that a reservoir computing exists that emulates given coupled maps by constructing a modularised network. We describe a possible mechanism for collapses of the emulation in the reservoir computing. Such transitory behaviour is caused by either (i) an escape from a finite-time stagnation near an unstable chaotic set, or (ii) a critical transition driven by the effective parameter drift. Our approach reveals the essential mechanism for reservoir computing and provides insights into the design of reservoir computing for practical applications.
@@ 1 Introduction
Machine learning, including the reservoir computing, has proven useful for model-free and data-driven predictions in a variety of scientific fields in recent years (see Lukoševičius and Jaeger (2009); Tanaka et al. (2019) and references therein). In studies on nonlinear phenomena, the reservoir computing is typically applied to large-scale phenomena such as chemical reaction Pathak et al. (2018), fluid turbulence Nakai and Saiki (2018); Kobayashi et al. (2021), and climate and environmental dynamics Paulin (2010); Mitsui and Boers (2021). The purpose of the reservoir computing is to emulate the dynamics $u(t+1)=g(u(t))$ behind time-series data $\{u(t)\}$ using the dynamics $\{r(t)\}$ that is generated by a large recurrent neural network $r(t+1)=f_{A}(r(t))$ with a network matrix $A$ Jaeger (2001); Jaeger and Haas (2004); Maass et al. (2002). Assuming that $f_{A}$ approximates any continuous functions by tuning $A$ when a sufficiently large network is used Cybenko (1989), the trivial solution to this problem is $r(0)=u(0),~{}f_{A}\simeq g$.
@@
As opposed to solving the non-convex optimisation problem to approximates $g$ with $f_{A}$ using a statistical optimisation method such as gradient descent, we may solve a simple convex optimisation problem using linear regression to obtain a projection of the network dynamics $P\{r(t)\}$, which is close to $\{u(t)\}$. The computation cost of the projection $P\{r(t)\}$ is clearly substantially smaller than that of the function approximation $f_{A}\simeq g$; however, we need to search for an appropriate initial network matrix $A$ for the successful emulation. The matrix $A$ is typically provided as a random matrix with a spectral radius $\rho(A)<1$ Jaeger (2001) (See Appendix A). One can search a number of the quenched random dynamics $\{r(t)\}$ by changing the random matrix $A$ to minimise $\lVert P\{r(t)\}-\{u(t)\}\rVert$. However, the design principle of such random matrices that satisfy appropriate quality, size, topology, and the order of the search cost is largely unknown.
@@
As a statistical interpretation, it is pointed out that the reservoir computing is some kind of nonlinear regression Gauthier et al. (2021). Indeed, a reservoir computing performed by a neural network without hidden layers $f_{O}$ is a simple perceptron, that is equivalent to logistic regression. It is also known that random neural networks also approximate $C^{1}$ functions arbitrary well Hart et al. (2020), which corresponds to the classical universal approximation theorem Cybenko (1989), and that the non-autonomous framework of reservoir computing is analogous to Takens’s embedding theorem, that works even under presence of noise Takens (2006); Sauer et al. (1991); Hart et al. (2020). See also Appendix B for the simplest example of a classical delay coordinate embedding performed by a non-autonomous reservoir computing.
@@
In this study, instead of a Monte Carlo search for the network matrix $A$ and a validation of the quality of the reservoir computing with randomly selected $A$, we construct the smallest possible network that can be included in $A$ to emulate the given dynamical systems. We analyse the behaviour of the constructed dynamical systems, including transitory behaviour that induces the collapse of the emulation Pathak et al. (2017); Kong et al. (2021). Our approach reveals the essential mechanism for reservoir computing and provides insights into the design of reservoir computing for practical applications.
@@ 2 Reservoir Computing
and expect the emulation $\{\bm{u}(t)\}\simeq\{\bm{y}(t)\}$.
@@ 3 Emulating coupled maps
which is equivalent to Eq. (3). Recall that at $t=0$ we obtain $x(0)=u(0)=y(0)$, and the dynamics is on $x=y=W_{\rm{out}}^{*}\bm{r}=v_{1}^{*}r^{(1)}+v_{2}^{*}r^{(2)}$, which may include the generalised synchronisation manifold in the reservoir $\cal{B}$ for the input $\{u(t)\}$ Lu et al. (2018); Lymburn et al. (2019). Thus, we achieve a perfect emulation $\{y(t)\}=\{u(t)\}$ for $t=0,1,2,\ldots$ by using the $K=2$ reservoir $\cal{B}$. We confirm that the above emulation is achieved in numerical experiments with linear regression (Fig. 2 (a)). Many other possible constructions are available for given one-dimensional maps.
@@
which is equivalent to Eq. (9). Noting that $\bm{x}(0)=\bm{u}(0)=\bm{y}(0)$, and the dynamics is on $\bm{x}=\bm{y}=W_{\rm{out}}^{*}\bm{r}$, we achieve a perfect emulation $\{\bm{y}(t)\}=\{\bm{u}(t)\}$.
We confirm that the above emulation is achieved in numerical experiments with linear regression (Fig. 2 (b)). Many other possible constructions are available for given coupled maps.
Figure 2: (a) Left:   map provided in Eq. (3) and return plots of $\{y(t)\}$ generated by constructed reservoir with $K=2$, and right:   attractor in $(r^{1)},r^{(2)},y)$, and the plane $y=W_{\rm{out}^{*}}\bm{r}$. (b) Map provided in Eq. (9) and return plots of $\{\bm{y}(t)\}$ generated by constructed reservoir with $K=4$. The parameters are $\gamma=0.9,~{}~{}\beta_{1}=0.75,~{}~{}\beta_{2}=0.25,~{}~{}\alpha_{1}=\alpha_{2}=0,~{}~{}v^{*}_{1}=-12,~{}~{}v_{2}^{*}=16,\epsilon=0.05$, and $\mu_{1}=\mu_{2}=\mu_{3}=\mu_{4}=0.01$.
In general, according to the property of neural networks as universal function approximators, we may emulate any one-dimensional maps
@@
where $h$ denotes an arbitrary continuous functions. For a large $K$, the projected dynamics $\{y(t)\}$ of a reservoir with $K$ neurons can be expressed as follows:
@@
For example, the projection $\{y(t)\}$ may emulate the logistic map $y(t+1)=a-y(t)^{2}$ when $h(y)\simeq-(1-\gamma)y+a-y^{2}$. Similarly, we can emulate arbitrary $M$-coupled maps with a linear coupling. To do that we use an $M$-dimensional projected dynamics of a reservoir with $KM$ neurons and a network matrix that consists of $M$ blocks of the size $K$. For example, with the construction in Eq. (9) where $f(y)\simeq a-y^{2}$, and the adequately optimised $W_{\rm{out}}^{*}$, the $M$-dimensional projected dynamics $\{\bm{y}(t)\}$ can emulate the dynamics of $M$-globally coupled logistic map $y^{(m)}(t+1)=(1-\epsilon)f(y^{(m)}(t))+\frac{\epsilon}{M}\sum_{i=1}^{M}f(y^{(i)}(t))$. This construction for coupled maps partially explains the successful emulation of PDEs by reservoir computing Pathak et al. (2017, 2018). Using a reservoir with a sufficiently large network, it is certainly possible to emulate the dynamics of a wide class of nonlinear PDEs, which are implemented on computers in the form of coupled maps. However, the essential problem is rather determining why a finite-size reservoir computing with a randomly selected network matrix efficiently emulates such dynamical systems with large degrees of freedom.
@@ 4 Emulation in large networks
While in gradient descent learning in multi-layer perceptrons, the degeneration of the network weights may cause vanishing gradient and decelerate learning Fukumizu and Amari (2000); Sato et al. (2022), the degeneration of the hidden layer dynamics in the reservoir computing may enhance precise emulations.
@@ 5 Collapse of emulation
$\eta(t+1)=(1-\gamma)\eta(t)+\gamma\sum_{k=1,2}q_{k}\tanh\left[\nu_{k}\eta(t)+\alpha_{k}\right].$ (27)
@@
When $\eta=0$, linear regression yields $W=W_{\rm{out}}^{*}$ and Eq. (26) is equivalent to Eq. (3). Thus, the collective excess neurons provides external perturbations to the projected dynamics $\{y(t)\}$. We present two examples of collapse of the emulation.
@@
In the first example, we set $\gamma=0.9,q_{1}=-12,q_{2}=16,\nu_{1}=0.75$, and $\nu_{2}=0.5$ satisfying $\rho(E)<1$. When $\alpha_{3}=\tanh^{-1}(2/3)$ and $\alpha_{4}=\tanh^{-1}(1/2)$, Eq. (27) has a tangency at the origin and a stable fixed point $\eta=\eta^{*}\simeq 3.773341$. In this case, the perfect emulation can be performed as a neutral solution on the centre manifold $\eta=0$. The chaotic saddle is approached when we start at an initial point in the effective basin defined by $\eta(0)<0$ and $\eta(0)\notin(-2.197527,-0.985279)$; otherwise, the dynamics converges to $\eta=\eta^{*}$ and the linear regression fails. If we are not at the onset of the bifurcation point but very close to it with a narrow channel, the dynamics of $\eta(t)$ remains near the origin for a long time. If the waiting time near the origin, which typically follows a power law, is longer than the training time $N$, we start at $\eta(0)\simeq\eta^{*}$ at time $t=0$ and the projected dynamics $\{y(t)\}$ emulates $\{u(t)\}$ for a finite time. However, in the long run, the dynamics of $\{\eta(t)\}$ escapes from the narrow channel near the origin and converges to the stable fixed point $\eta=\eta^{*}$, and the emulation collapses (FIG 3 (a)).
Figure 3: (a) Return plots and dynamics of the emulation collapse caused by a tangency. The parameters are $q_{1}=-12,q_{2}=16,\nu_{1}=0.75,\nu_{2}=0.5,\alpha_{3}=\tanh^{-1}(2/3)$, and $\alpha_{4}=\tanh^{-1}(1/2)$. The black line indicates Eqs. (26) and (27). The origin is neutrally stable and the fixed point $\eta=\eta^{*}\simeq 3.773341$ is stable in Eq. (27). (b) Return plots and dynamics of the emulation collapse caused by a critical transition. The parameters are $q_{1}=1,\nu_{1}=1-10^{-8}$, and $\alpha_{3}=0.001$. The black line indicates Eqs. (26) and (28). A narrow channel near the origin and a stable fixed point $\eta=\eta^{*}\simeq 0.143626$ exist.
@@
The second example is based on only one excess neuron. We set $\gamma=0.9,q_{1}=1,\nu_{1}=1-10^{-8}$, and $\alpha_{3}=0.001$ satisfying $\rho(E)<1$ and obtain a hyperbolic tangent map
$\eta(t+1)=(1-\gamma)\eta(t)+\gamma\tanh\left[\nu_{1}\eta(t)+\alpha_{3}\right].$ (28)
@@
In this case, the perfect emulation can be performed on an unstable chaotic set. The dynamics of Eq. (28) does not have tangencies but may have a narrow channel near the origin. It has a stable fixed point $\eta=\eta^{*}\simeq 0.143626$. The origin can be approached when we start at $\eta(0)<0$; otherwise, the dynamics converges to the stable fixed point $\eta=\eta^{*}$ and the linear regression fails. A very slow uniform motion $\eta(t)\simeq\gamma\tanh(\alpha_{3})t\simeq 10^{-4}t$ occurs near the origin, which functions as a parameter drift in Eq. (26). Thus, considering that other excess neurons function as external noise in a large network, bifurcation that is similar to the critical transition may occur Scheffer et al. (2009); Kong et al. (2021). We observe orbits that are close to the training data for a finite time. However, before $\eta$ reaches to the stable fixed point $\eta=\eta^{*}$, it may arrive at the tipping point of Eq. (26) and the emulation collapses (FIG 3 (b)). In these two scenarios, the initial matrices and the initial conditions of $\bm{r}(0)$ have a positive measures in the parameter and state space. Thus, the above phenomena can be observed in a quenched random dynamical systems with a random matrix $A$ and with an initial condition $\bm{r}(0)$.

### v3

@@ Abstract
A minimal model for reservoir computing is studied. We demonstrate that a reservoir computer exists that emulates given coupled maps by constructing a modularized network. We describe a possible mechanism for collapses of the emulation in the reservoir computing by introducing a measure of finite scale deviation. Such transitory behaviour is caused by either (i) an escape from a finite-time stagnation near an unstable chaotic set, or (ii) a critical transition driven by the effective parameter drift. Our approach reveals the essential mechanism for reservoir computing and provides insights into the design of reservoir computer for practical applications.
@@ 1 Introduction
Machine learning, including the reservoir computing, has proven useful for model-free and data-driven predictions in a variety of scientific fields in recent years (see [1, 2] and references therein). In studies on nonlinear phenomena, the reservoir computing is typically applied to large-scale phenomena such as chemical reaction [3], fluid turbulence [4, 5], and climate and environmental dynamics [6, 7]. The purpose of the reservoir computing is to emulate the dynamics $u(t+1)=g(u(t))$ behind time-series data $\{u(t)\}$ using the dynamics $\{r(t)\}$ that is generated by a large recurrent neural network $r(t+1)=f_{A}(r(t))$ with a network matrix $A$ [8, 9, 10]. Assuming that $f_{A}$ approximates any continuous functions by tuning $A$ when a sufficiently large network is used [11], the trivial solution of this problem is $r(0)=u(0),~{}f_{A}=g$.
@@
As opposed to solving the non-convex optimisation problem to approximates $g$ with $f_{A}$ using a statistical optimisation method such as gradient descent, we may solve a simple convex optimisation problem using linear regression to obtain a projection of the network dynamics $P\{r(t)\}$, which is close to $\{u(t)\}$. The computation cost of the projection $P\{r(t)\}$ is clearly substantially smaller than that of the function approximation of $g$; however, we need to search for an appropriate initial network matrix $A$ for the successful emulation. The matrix $A$ is typically provided as a random matrix with a spectral radius $\rho(A)<1$ [8] (See Appendix A). One can search a number of the quenched random dynamics $\{r(t)\}$ by changing the random matrix $A$ to minimise $\lVert P\{r(t)\}-\{u(t)\}\rVert$. However, the design principle of such random matrices that satisfy appropriate quality, size, topology, and the order of the search cost is largely unknown.
@@
As a statistical interpretation, it is pointed out that the reservoir computing is some kind of nonlinear regression [12]. Indeed, a reservoir computing performed by a neural network without hidden layers $f_{O}$ is a simple perceptron, that is equivalent to logistic regression. It is also known that random neural networks also approximate $C^{1}$ functions arbitrary well [13], which corresponds to the classical universal approximation theorem [11], and that the non-autonomous framework of reservoir computing is analogous to Takens’s embedding theorem, that works even under presence of noise [14, 15, 13]. See also Appendix B for the simplest example of a classical delay coordinate embedding performed by a non-autonomous reservoir computer.
@@
In this study, instead of a Monte Carlo search for the network matrix $A$ and a validation of the quality of the reservoir computing with randomly selected $A$, we construct the smallest possible network that can be included in $A$ to emulate the given dynamical systems. We analyse the behaviour of the constructed dynamical systems, including transitory behaviour that induces the collapse of the emulation [16, 17]. Our approach reveals the essential mechanism for reservoir computing and provides insights into the design of reservoir computer for practical applications.
@@ 2 Reservoir Computing
and expect the emulation $\{\bm{u}(t)\}\approx\{\bm{y}(t)\}$.
@@ 3 Emulating coupled maps
which is equivalent to Eq. (3). Recall that at $t=0$ we obtain $x(0)=u(0)=y(0)$, and the dynamics is on $x=y=W_{\rm{out}}^{*}\bm{r}=v_{1}^{*}r^{(1)}+v_{2}^{*}r^{(2)}$, which may include the generalised synchronisation manifold in the reservoir $\cal{B}$ for the input $\{u(t)\}$ [18, 19, 20]. Thus, we achieve a perfect emulation $\{y(t)\}=\{u(t)\}$ for $t=0,1,2,\ldots$ by using the $K=2$ reservoir $\cal{B}$. We confirm that the above emulation is achieved in numerical experiments with linear regression (Fig. 2 (a)). Many other possible constructions are available for given one-dimensional maps.
Figure 2: (a) Left:   the map provided in Eq. (3) and return plots of $\{y(t)\}$ generated by constructed reservoir with $K=2$, and right:   the attractor in $(r^{1)},r^{(2)},y)$, and the plane $y=W_{\rm{out}^{*}}\bm{r}$. (b) The map provided in Eq. (9) and return plots of $\{\bm{y}(t)\}$ generated by constructed reservoir with $K=4$. The parameters are $\gamma=0.9,~{}~{}\beta_{1}=0.75,~{}~{}\beta_{2}=0.25,~{}~{}\alpha_{1}=\alpha_{2}=0,~{}~{}v^{*}_{1}=-12,~{}~{}v_{2}^{*}=16,\epsilon=0.05$, and $\mu_{1}=\mu_{2}=\mu_{3}=\mu_{4}=0.01$.
@@
which is equivalent to Eq. (9). Noting that $\bm{x}(0)=\bm{u}(0)=\bm{y}(0)$, and the dynamics is on $\bm{x}=\bm{y}=W_{\rm{out}}^{*}\bm{r}$, we achieve a perfect emulation $\{\bm{y}(t)\}=\{\bm{u}(t)\}$. We confirm that the above emulation is achieved in numerical experiments with linear regression (Fig. 2 (b)).
In general, according to the property of neural networks as universal function approximators, with sufficiently many neurons, we may emulate any one-dimensional maps
@@
where $h$ denotes an arbitrary continuous functions. In practice, it is impossible to learn structurally unstable dynamical systems with finite data, however, in machine learning, input data always includes observational noise and/or finite size fluctuation, and thus function approximabilities is typically discussed in theory. For a large $K$, the projected dynamics $\{y(t)\}$ of a reservoir with $K$ neurons can be expressed as follows:
@@
For example, the projection $\{y(t)\}$ may emulate the logistic map $y(t+1)=a-y(t)^{2}$ when $h(y)=-(1-\gamma)y+a-y^{2}$. Similarly, we can emulate arbitrary $M$-coupled maps with a linear coupling. To do that we use an $M$-dimensional projected dynamics of a reservoir with $KM$ neurons and a network matrix that consists of $M$ blocks of the size $K$. For example, with the construction in Eq. (9) where $f(y)=a-y^{2}$, and the adequately optimised $W_{\rm{out}}^{*}$, the $M$-dimensional projected dynamics $\{\bm{y}(t)\}$ can emulate the dynamics of $M$-globally coupled logistic map $y^{(m)}(t+1)=(1-\epsilon)f(y^{(m)}(t))+\frac{\epsilon}{M}\sum_{i=1}^{M}f(y^{(i)}(t))$. This construction for coupled maps partially explains the successful emulation of PDEs by reservoir computers [16, 3]. Using a reservoir with a sufficiently large network, it is certainly possible to emulate the dynamics of a wide class of nonlinear PDEs, which are implemented on computers in the form of coupled maps. However, the essential problem is rather why a finite-size reservoir computer with a randomly selected network matrix efficiently emulates such dynamical systems with large degrees of freedom.
@@ 4 Emulation in large networks
While in gradient descent learning in multi-layer perceptrons, the degeneration of the network weights may cause vanishing gradient and decelerate learning process [21, 22], the degeneration of the hidden layer dynamics in the reservoir computing may enhance precise emulations.
5 Effectively successful emulation
To evaluate emulation of one-dimensional maps $u(t+1)=g(u(t))\in[s_{1},s_{2}]$, we introduce a finite scale deviation $D_{t_{0}}(T,\delta)$ in a time window $t_{0}-T\leq t\leq t_{0}+T$ between the input data $\{u(t)\}$ generated by the dynamical system $g$, and the emulation $\{y(t)\}$ by a reservoir computer, based on an observation within a finite time $2T$ and a finite spatial resolution $\delta$. First, we set a partition $\{\Delta_{i}\}$ with $m$ bins of the size $\delta$ for the interval $[s_{1},s_{2}]$ as $\Delta_{i}=[s_{1}+(i-1)\delta,s_{1}+i\delta],~{}~{}(i=1,\ldots,m,~{}m\delta=|s_{2}-s_{1}|)$, and observe the dynamics $\{y(t)\}$ visiting the $i$th small interval $\Delta_{i}$. When we observe the dynamics in $\Delta_{i}$ for $k^{(i)}>0$ times at $t=\tau_{1}^{(i)},\ldots,\tau_{k}^{(i)}$, we take the center of mass of the tuple $(y(\tau^{(i)}),y(\tau^{(i)}+1))$ over the generated data, given as
$(\tilde{y}^{(i)},\tilde{y}^{(i)}{}^{\prime})=\left(\frac{y(\tau_{1}^{(i)})+\cdots y(\tau_{k^{(i)}}^{(i)})}{k^{(i)}},~{}\frac{y(\tau_{1}^{(i)}+1)+\cdots y(\tau_{k^{(i)}}^{(i)}+1)}{k^{(i)}}\right),$ (23)
where $y(\tau_{1}^{(i)}),\ldots,y(\tau_{k^{(i)}}^{(i)})\in\Delta_{i}$. The tuple $(\tilde{y}^{(i)},\tilde{y}^{(i)}{}^{\prime})$ is the representative of successive observations of the dynamics in $\Delta_{i}$. When the dynamics does not visit $\Delta_{i}$, that is $k^{(i)}=0$, we define $(\tilde{y}^{(i)},\tilde{y}^{(i)}{}^{\prime})=(s_{1}+(i-\frac{1}{2})\delta,0)$. Then the finite scale deviation measured in a time window $t_{0}-T\leq t\leq t_{0}+T$ is defined as
$D_{t_{0}}(T,\delta)=\frac{1}{m}\sum_{i=1}^{m}(g(\tilde{y}^{(i)})-\tilde{y}^{(i)}{}^{\prime})^{2},$ (24)
where the limit $D_{t_{0}}^{\infty}=\lim_{T\to\infty}\lim_{\delta\to 0}D_{t_{0}}(T,\delta)$, approximating the deviation between $\{(\tilde{y}^{(i)},g(\tilde{y}^{(i)}))\}$ and $\{(\tilde{y}^{(i)},\tilde{y}^{(i)}{}^{\prime})~{}~{}(i=1,\ldots,m)\}$ in a given finite scale. For the perfect emulation, we have $D_{t}^{\infty}=0$ for all $t$. We may estimate $g(\tilde{y}^{(i)})$ with the input data by a linear interpolation in case that we do not know $g$. If a criterion $D_{t_{0}}(T,\delta)<d$ is satisfied, we say that the simulation is effectively successful, in a time window $t_{0}-T\leq t\leq t_{0}+T$, with a deviation $d$, in a finite time $2T$, and in a finite spatial resolution $\delta$.
As an example, we compute the deviation $D_{t_{0}}(T,\delta)$ with $t_{0}=1500$, $T=1500$, $\delta=10^{-3}$, and $d=5\times 10^{-4}$ for the reservoir computing given in Eqs. (3)-(6). The parameters are given as $\mu_{1}=\mu_{2}=0.01$, $\beta_{1}=0.75,\beta_{2}=0.25$, $\alpha_{1}=\alpha_{2}=0$, and the matrix $B$ is given as the following random matrix;
$B=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\\ \end{pmatrix}=\begin{pmatrix}-0.12&0.16\\ -0.12&0.16\\ \end{pmatrix}+\begin{pmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\\ \end{pmatrix},$ (25)
where $\xi_{ij}\in[-0.8,0.8]$ randomly chosen from uniform distribution. In FIG 3, the black points plotted near a section $(b_{21},b_{22})=(-0.12,0.16)$, indicate the effectively successful emulation after $N=10^{4}$ training phase with a criterion $D_{t_{0}}(T,\delta)<d$. The red point indicates the perfect emulation, with $D_{t}^{\infty}=0$ for all $t$, given by $(b_{11},b_{12})=(-1.2,1.6)$, which induces the correct estimation of the original model with $(v_{1}^{*},v_{2}^{*})=(-12,16)$. Around the perfect emulation point, there is a region consists of the black points corresponding to effectively successful emulation.
Figure 3: The perfect emulation surrounded by effectively successful emulation: The black points plotted near a section $(b_{21},b_{22})=(-0.12,0.16)$, indicate the effectively successful emulation after $N=10^{4}$ training phase with a criterion $D_{t_{0}}(T,\delta)<d$. The red point indicates the perfect emulation given by $(b_{11},b_{12})=(-1.2,1.6)$. Around the perfect emulation point, there is a region consists of the black points corresponding to effectively successful emulation. The parameters are given as $t_{0}=1500$, $T=1500$, $\delta=10^{-3}$, and $d=5\times 10^{-4}$.
Note that the region consists of the black dots in FIG 3 includes ”nearly” successful emulation achieved by transient chaos, due to our criterion based on the finite space-time deviation. In the region far from the perfect emulation point, the estimated dynamical system is possibly after a bifurcation to a periodic window. In such cases, although the dynamics in the autonomous prediction phase starts with $y(0)=u(0)$, where $u(0)$ is on the attractor of the original dynamics $\{u(t)\}$, the initial point $y(0)$ is not on the original attractor, but is very close to the chaotic saddle of the projected dynamics $\{y(t)\}$ and may show a very long transient chaos similar to the original dynamics.
@@ 6 Collapse of emulation
$\eta(t+1)=(1-\gamma)\eta(t)+\gamma\left(q_{1}\tanh\left[\nu_{1}\eta(t)+\alpha_{3}\right]+q_{2}\tanh\left[\nu_{2}\eta(t)+\alpha_{4}\right]\right).$ (30)
@@
When $\eta(t)=0$, linear regression yields $W=W_{\rm{out}}^{*}$ and Eq. (29) is equivalent to Eq. (3). Thus, the collectives of excess neurons provides external perturbations to the projected dynamics $\{y(t)\}$. We present two examples of the collapse of the emulation in these types.
@@
In the first example, we set $\gamma=0.9,q_{1}=-12,q_{2}=16,\beta_{1}=\nu_{1}=0.75$, $\beta_{2}=0.25$, $\nu_{2}=0.5$, and $\alpha_{1}=\alpha_{2}=0$. When $\alpha_{3}=\tanh^{-1}(2/3)$ and $\alpha_{4}=\tanh^{-1}(1/2)$, Eq. (30) has a tangency at the origin and a stable fixed point $\eta=\eta^{*}\simeq 3.773542$. In this case, the perfect emulation can be performed as a neutral solution on the centre manifold $\eta(t)=0$. The chaotic saddle is approached when we start at an initial point in the effective basin defined by $\eta(0)<0$ and $\eta(0)\notin(-2.197527,-0.985279)$; otherwise, the dynamics converges to $\eta(t)=\eta^{*}$ and the emulation fails. In general, if the system is at the onset of a bifurcation and have a narrow channel at the stagnation point, the dynamics stays near the stagnation point for a long time. The duration time near the stagnation point typically follows a power law, and can be arbitrarily longer. If the dynamics of $\eta(t)$ is near the stagnation point, and the resulting waiting time is longer than the training time $N$, the projected dynamics $\{y(t)\}$ can emulate $\{u(t)\}$ for a finite time. In particular, when $\eta(t)=0$, the perfect emulation is achieved. However, in the long run, the dynamics of $\eta(t)$ escapes from the neighbourhood of the stagnation point and converges to the stable fixed point $\eta(t)=\eta^{*}\simeq 3.773542$. Thus, the emulation is effectively successful for a finite time but eventually collapses (FIG 4).
Figure 4: The emulation collapse caused by a tangency: The origin is almost neutrally stable, and the fixed point $\eta(t)=\eta^{*}\simeq 3.773542$ is stable in Eq. (30). The dynamics of $\eta(t)$ and the deviation $D_{t}(T,\delta)$, where $T=1500,~{}\delta=10^{-3}$, is depicted in the blue and the black line, respectively, in the upper panel. The red part of $D_{t}$ is below the threshold $D_{t}<d=0.0001$. In the lower panel, the time series of $y(t)$ is depicted in black dots, and, in the red part, the emulation is effectively successful. The inset of the upper panel is the return plot of $\{y(t)\}$, and the red part is the finite time return plot of $\{y(t)\}$ during effectively successful emulation. The parameters are $\gamma=0.9$, $q_{1}=-12,q_{2}=16,\beta_{1}=\nu_{1}=0.75,\beta_{2}=0.25,\nu_{2}=0.5,\alpha_{1}=\alpha_{2}=0,\alpha_{3}=\tanh^{-1}(2/3)-10^{-8}$, and $\alpha_{4}=\tanh^{-1}(1/2)$.
Figure 5: The emulation collapse caused by a parameter drift: There is an almost neutrally stable narrow channel near the origin, and the fixed point $\eta(t)=\eta^{*}\simeq 0.019036$ is stable in Eq. (31). The dynamics of $\eta(t)$ and the deviation $D_{t}(T,\delta)$, where $T=1500,~{}\delta=10^{-3}$ is depicted in the blue and the black line, respectively, in the upper panel. The red part of $D_{t}$ is below the threshold $D_{t}<d=0.005$. In the lower panel, the time series of $y(t)$ is depicted in black dots, and in the red part, the emulation is effectively successful. The inset of the upper panel is the return plot of $\{y(t)\}$, and the red part is the finite time return plot of $\{y(t)\}$ during effectively successful emulation. The parameters are $\gamma=0.9$, $q_{1}=1,\nu_{1}=1$, and $\alpha_{3}=2.3\times 10^{-6}$.
@@
The second example is based on only one excess neuron with $\eta(t+1)=q_{1}r^{(3)}(t)$. We set $\gamma=0.9,q_{1}=1,\nu_{1}=1$, and $\alpha_{3}=2.3\times 10^{-6}$, and obtain a hyperbolic tangent map
$\eta(t+1)=(1-\gamma)\eta(t)+\gamma q_{1}\tanh\left[\nu_{1}\eta(t)+\alpha_{3}\right],$ (31)
@@
where $\eta(0)=-0.03$. In this case, the perfect emulation can be performed on an unstable chaotic set. The dynamics of Eq. (31) does not have a tangency but has an almost neutrally stable narrow channel near the origin, and has a stable fixed point $\eta(t)=\eta^{*}\simeq 0.019036$. The origin can be approached when we start at $\eta(0)<0$; otherwise, the dynamics converges to the stable fixed point $\eta(t)=\eta^{*}$ and the emulation fails. A very slow uniform motion $\eta(t)\simeq\gamma\tanh(\alpha_{3})t\simeq 10^{-6}t$ emerges near the origin, which works as a slow parameter drift in Eq. (29). Thus, considering a situation that other neurons in the large network work as external noise, the critical transition may occur with the parameter drift and the external noise [23, 17]. We observe orbits that are close to the original dynamics for a finite time, however, before $\eta$ reaches to the stable fixed point $\eta(t)=\eta^{*}\simeq 0.019036$, it may arrive at the tipping point of Eq. (29) and the emulation collapses by the critical transition (FIG 5).
In these two scenarios, the initial matrices and the initial conditions of $\bm{r}(0)$ have a positive measures in the parameter and state space. Thus, the above phenomena can be observed in a quenched random dynamical systems with a random matrix $A$ and with an initial condition $\bm{r}(0)$.